If a=i-j and b=i+k then a unit vector coplanar with a and b and perpendicular to a is

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

you can take any vector in plane of two given vector

r =xa +yb

 =x(i-j) +y (i+k)

 =i(x+y) -xj +yk

 

if it is perpendicular to a then

 r.a=o

{i(x+y) -xj +yk}.(i-j)=0

2x+y=0

so r will become r= -x[i+j +2k]

 if it is a unit vecttor then ( x2 + x2 +4x2)1/2=1

                                       x= ±1/√6

so r= ±1/√6[i +j+2k]


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin